Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 374     Entry time: 29 Sep 2023, 22:59
Author: Laurent Mercadier 
Group: SCS 
Subject: Shift summary 

# Late shift summary 29.09.2023

Laurent, Phuong
## Achievements
- Continue tr-RIXS
- Laser fluence scan
- Check PPL 266 nm power
- Setup DAMNIT

## Problems:
- online analysis not displayed on gui1 server fixed by DOC

## Plan for the night:
- continue tr-RIXS
- monitor energy and position drifts

ELOG V3.1.4-7c3fd00